OpenStudy (anonymous):

4. Let f(x) = 3x3 – 4x – 1 and g(x) = x + 1. Find 5. Let f(x) = x4 – 8x3 + 16x2 – 19 and g(x) = x – 5. Find 6. Create your own third degree polynomial that when divided by x + 2 has a remainder of –4. 7. Create your own division of polynomials problem. Demonstrate how this problem would be solved using both long division and synthetic division.

OpenStudy (anonymous):

wats ur question ?????

OpenStudy (anonymous):

and for # 4 & 5 were looking to find F(X)/G(X)

OpenStudy (anonymous):

4. \[F(x)=\int\limits\limits_{}^{}f(x)dx\] \[F(x)=(3/4)x ^{4}-2x ^{2}-x+C\] \[G(x)=\int\limits\limits\limits_{}^{}g(x)dx\] \[G(x)=(1/5)x^{5}-2x^{4}-(16/3)x^{3}-19x+C\] \[F(x)/G(x)=((3/4)x ^{4}-2x ^{2}-x+C)/((1/5)x^{5}-2x^{4}-(16/3)x^{3}-19x+C)\]

OpenStudy (anonymous):

you can solve #5 similarly

OpenStudy (anonymous):

6. \[(2x ^{3}+2x ^{2}+2x-8)\div(x/2)\]has a remainder of -4 if you didn't mean to say "divided by x + 2" and you meant to say "divided by x-2", then \[x^3+2 x^2-5 x-10\] when divided by "x-2" has a remainder of -4